OpenStudy (anonymous):

Simplify: 14x - 14 - 2 {x-[3x +(13x-3) ] } + 28 please show work

OpenStudy (anonymous):

First work with the insides. Thus we should work with 3x+(13x-3) first. This gives 16x - 3. Substituting what we've got back to the whole equation, it now looks like this: 14x - 14 - 2 {x-(16x-3) } + 28 Now lets work with the insides again by simplifying (1)x-(16x-3), which should give the following: -15x-3. Substituting what we've got back to the whole equation, it now looks like this: 14x - 14 - 2 (-15x -3) + 28 Now we need to yet again work with the insides: -2 (-15x -3) = 30x +6 We substitute this back into the equation and get: 14x - 14 +30x +6 + 28 Now simplifying the equation further, we get the following after some simple algebra: 44x + 8

OpenStudy (anonymous):

=14x-16{x-[3x+13x-3]=14x-16{x-16x-3}=14x-16x+256x+48=254x+48

OpenStudy (anonymous):

kwo is wrong x-(16x-3) -> -15x + 3

OpenStudy (anonymous):

TheMind why -16 { }.. it is -2 { }

OpenStudy (anonymous):

ooppsss.... yes, Coolsector is right about that

OpenStudy (anonymous):

yes a little misspell i guess

OpenStudy (anonymous):

** modified and corrected ** First work with the insides. Thus we should work with 3x+(13x-3) first. This gives 16x - 3. Substituting what we've got back to the whole equation, it now looks like this: 14x - 14 - 2 {x-(16x-3) } + 28 Now lets work with the insides again by simplifying (1)x-(16x-3), which should give the following: -15x+3. Substituting what we've got back to the whole equation, it now looks like this: 14x - 14 - 2 (-15x + 3) + 28 Now we need to yet again work with the insides: -2 (-15x +3) = 30x -6 We substitute this back into the equation and get: 14x - 14 +30x -6 + 28 Now simplifying the equation further, we get the following after some simple algebra: 44x + 8

OpenStudy (anonymous):

44x+8

OpenStudy (anonymous):

44x+8 is the final answer?

OpenStudy (anonymous):

Yes, and all the work is shown above in my previous comment. However, if you want to solve for x, then it would be -8/44, which gives -2/11 or -0.18181818

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

I am assuming, however, that 0 = 44x + 8 initially. Just so you know...